A couple plans to have three children. What is the probability of getting at least one girl? 7/8 1/8 1/2 3/4
Answer by venugopalramana(3286)

A couple plans to have three children. What is the probability of getting at least one girl? 7/8 1/8 1/2 3/4
PGIRL=PBOY=0.5
PROBABILITY OF NO GIRL IN 3 =(1/2)*(1/2)*(1/2)=1/8
HENCE PROBABILITY OF HAVING ATLEAST ONE GIRL =1-1/8=7/8